# Coordinated Multi-Armed Bandits for Improved Spatial Reuse in Wi-Fi

### v1

@@ I Introduction
This paper investigates the potential of MAPC-enabled coordinated Multi-Armed Bandits (MABs) for enhancing SR in coordinated Wi-Fi networks. Our approach presents a promising alternative to C-SR, with the aim of achieving similar network-wide benefits while reducing the complexity of joint Transmit Power Control (TPC) and PD adjustment decisions. MAB is a popular framework for sequential decision-making under uncertainty, which makes it well-suited for overcoming the complex OBSS interactions that occur in Wireless Local Area Networks (WLANs). This underlying complexity is precisely the main motivation for adopting Artificial Intelligence (AI) and Machine Learning (ML) techniques such as MAB, which are expected to address the dynamic and unpredictable nature of wireless environments, for which static approaches (even if coordinated) can fail. In fact, the adoption of AI/ML in the IEEE 802.11 is gaining momentum with the establishment of the AI/ML Study Group (SG) and the AI/ML Standing Committee (SC) [5]. This paper’s contributions are as follows:
@@ III Coordinated MAB
We consider a Multi-Agent MAB (MA-MAB) setting in which multiple agents are individually enrolled in different BSSs for optimizing SR by tuning the PD (which determines to which extent inter-BSS interference can be ignored) and the transmit power (which determines to which extent inter-BSS interference is created). Moreover, by leveraging MAPC capabilities, we allow agents to exchange feedback about their performance regularly, thus aiming at creating collaborative strategies that enhance global performance. Formally, based on some action–selection strategy $\mathcal{E}$ (cf. Section III-A), each player (or agent) $p\in\mathcal{P}$ sequentially plays an action (or arm) $k\in\mathcal{A}^{(p)}$ at each time step $t=1,...,T$ and obtains a reward $r^{(p)}(t+1)\in\mathbb{R}$ that depends on the joint action profile $K(t)=\{k^{(1)}(t),...,k^{(|\mathcal{P}|)}(t)\}$ and on the reward calculation strategy $\mathcal{R}$ (cf. Section III-B). For the SR problem, we consider actions composed of discrete PD ($\gamma$) and transmit power ($\zeta$) values, which leads to an individual action space $\mathcal{A}^{(p)}_{|\gamma|\times|\zeta|}$ and a global action space $\mathcal{A}_{|\mathcal{P}|\times|\gamma|\times|\zeta|}$.
Figure 1: Representation of an agent’s operation in an OBSS, where the rewards associated with the different available actions are drawn from an unknown distribution that stems from the joint SR configuration.
@@ IV Performance Evaluation
In this section, we study the different strategies for improving SR through simulations. We use Komondor [17], an IEEE 802.11 simulator with embedded AI-capable agents. The various MAB implementations were simulated along with the network operation using the parameters collected in Table I. More details on the considered IEEE 802.11 frame types, sizes, and interframe periods are in [4, Table B.6].
@@ IV-A Interactions between 2 BSSs
Figure 2: Considered 2-BSS deployment (toy scenario) and average performance achieved when using different power levels and PD configurations ($A_{1}=\{10,-72\}$ dBm, $A_{2}=\{10,-82\}$ dBm, $A_{3}=\{20,-72\}$ dBm, $A_{4}=\{20,-82\}$ dBm).
@@
Figure 3: Average throughput observed in the toy scenario when applying OBSS/PD SR, uncoordinated bandit (Uncoord-Bandit) with $\varepsilon$-greedy, and coordinated bandit (Coord-Bandit) with $\varepsilon$-greedy (AVG).
Table II: Percentage of times each action is selected by uncoordinated bandit (Uncoord-Bandit) and coordinated bandit (Coord-Bandit) in the toy scenario.
@@ IV-B Random scenarios
Figure 6: Considered 9 BSS deployment with randomly located stations.
@@
Figure 7: Mean performance achieved in the $9$ BSS random deployment for the different considered coordinated MAB implementations ($\varepsilon$-greedy and Thompson sampling) and shared reward strategies (AVG, MAX-MIN, and PF). Left: mean throughput, left-center: mean average delay, right-center: mean airtime, right: percentage of time in NAV.
@@
Based on previous observations, we focus on the performance achieved by the best-performing options for coordinated bandits, namely $\varepsilon$-greedy (AVG) and Thompson sampling (PF). Figure 8 shows the Cumulative Distribution Function (CDF) of the mean throughput and channel access delay experienced throughout all the simulations.
@@
The two MAB implementations grant higher throughput than OBSS/PD SR, $\varepsilon$-greedy (AVG) being the best one (Figure 8). In some cases, OBSS/PD SR leads to slightly higher throughput than Thompson sampling (PF), but this stems from the unfair situations observed above. For mean access delay, the two MAB implementations lead to similar results, but in both cases OBSS/PD SR is substantially outperformed.
@@ V Conclusions
This paper explored the potential of coordinated MA-MAB to improve SR in MAPC-equipped Wi-Fi networks. Our findings demonstrate that coordinated MABs offer a significant advantage over the OBSS/PD SR operation and uncoordinated MABs. Moreover, we studied different exploration-exploitation strategies and showed that explicit exploration ($\varepsilon$-greedy) grants control on the exploration but might not adapt properly to different situations, whereas implicit exploration (Thompson sampling) leads to higher instability but adapts well to various contexts and types of rewards and eventually leads to optimal policies. In addition, we studied different reward-sharing strategies and showed that metrics such as AVG and PF can drive agents to find global configurations that maximize overall performance. Meanwhile MAX-MIN is useful to maximize minimum performance, but fails to maximize overall performance in some situations.

### v2

@@ I Introduction
This paper investigates the potential of MAPC-enabled coordinated Multi-Armed Bandits (MABs) for enhancing SR in coordinated Wi-Fi networks. Our approach presents a promising alternative to C-SR, with the aim of achieving similar network-wide benefits while reducing the complexity of joint Transmit Power Control (TPC) and PD adjustment decisions. MAB is a popular framework for sequential decision-making under uncertainty, which makes it well-suited for overcoming the complex OBSS interactions that occur in Wireless Local Area Networks (WLANs). This underlying complexity is precisely the main motivation for adopting Artificial Intelligence (AI) and Machine Learning (ML) techniques such as MABs, which are expected to address the dynamic and unpredictable nature of wireless environments, for which static approaches (even if coordinated) can fail. In fact, the adoption of AI/ML in the IEEE 802.11 is gaining momentum with the establishment of the AI/ML Study Group (SG) and the AI/ML Standing Committee (SC) [5]. This paper’s contributions are as follows:
@@ III Coordinated MAB
We consider a Multi-Agent MAB (MA-MAB) setting in which multiple agents are individually enrolled in different BSSs for optimizing SR by tuning the PD (which determines to which extent inter-BSS interference can be ignored) and the transmit power (which determines to which extent inter-BSS interference is created). The proposed solution departs from the operations defined for OBSS/PD-based SR, which was detailed in [4], hence making it compatible with currently available products. As a step further, and by leveraging MAPC capabilities, we allow agents to exchange feedback about their performance regularly, thus aiming at creating collaborative strategies that enhance global performance. Formally, based on some action–selection strategy $\mathcal{E}$ (cf. Section III-A), each player (or agent) $p\in\mathcal{P}$ sequentially plays an action (or arm) $k\in\mathcal{A}^{(p)}$ at each time step $t=1,...,T$ and obtains a reward $r^{(p)}(t+1)\in\mathbb{R}$ that depends on the joint action profile $K(t)=\{k^{(1)}(t),...,k^{(|\mathcal{P}|)}(t)\}$ and on the reward calculation strategy $\mathcal{R}$ (cf. Section III-B). For the SR problem, we consider actions composed of discrete PD ($\gamma$) and transmit power ($\zeta$) values, which leads to an individual action space $\mathcal{A}^{(p)}_{|\gamma|\times|\zeta|}$ and a global action space $\mathcal{A}_{|\mathcal{P}|\times|\gamma|\times|\zeta|}$.
Figure 1: Representation of an agent’s operation in an OBSS, where the rewards associated with the different available actions are drawn from an unknown distribution.
@@ IV Performance Evaluation
In this section, we study the different strategies for improving SR through simulations. We use Komondor [17], an IEEE 802.11 simulator with embedded AI-capable agents. The various MAB implementations were simulated along with the network operation using the parameters collected in Table I. More details on the considered IEEE 802.11 frame types, sizes, and inter-frame periods are in [4, Table B.6].
@@ IV-A Interactions between 2 BSSs
Figure 2: Considered 2-BSS toy deployment and mean performance of $A_{1}=\{10,-72\}$ dBm, $A_{2}=\{10,-82\}$ dBm, $A_{3}=\{20,-72\}$ dBm, $A_{4}=\{20,-82\}$ dBm.
@@
Figure 3: Average throughput observed in the toy scenario for OBSS/PD SR, uncoordinated bandit with $\varepsilon$-greedy, and coordinated bandit with $\varepsilon$-greedy (AVG).
Table II: Selected actions by uncoordinated (Uncoord-Bandit) and coordinated bandits (Coord-Bandit) in the toy scenario.
@@ IV-B Random scenarios
Figure 6: Considered 9-BSS random deployment.
@@
[…]
@@
Based on previous observations, we focus on the performance achieved by the best-performing options for coordinated bandits, namely $\varepsilon$-greedy (AVG) and Thompson sampling (PF). Figure 8 shows the Cumulative Distribution Function (CDF) of the mean throughput and channel access delay experienced throughout all the simulations. The two MAB implementations grant higher throughput than OBSS/PD SR, $\varepsilon$-greedy (AVG) being the best one (Figure 8). In some cases, OBSS/PD SR leads to slightly higher throughput than Thompson sampling (PF), but this stems from the unfair situations observed above. For mean access delay, the two MAB implementations lead to similar results, but in both cases OBSS/PD SR is substantially outperformed.
@@ V Conclusions
This paper explored the potential of coordinated MA-MAB to improve SR in MAPC-equipped Wi-Fi networks. Our findings demonstrate that coordinated MABs offer a significant advantage over the OBSS/PD SR operation and uncoordinated MABs. We studied different exploration-exploitation strategies and showed that explicit exploration ($\varepsilon$-greedy) grants control on the exploration but might not adapt properly to different situations, whereas implicit exploration (Thompson sampling) leads to higher instability but adapts well to various contexts and reward types, eventually leading to optimal policies. In addition, we studied different reward-sharing strategies and showed that metrics such as AVG and PF can drive agents to find global configurations that maximize overall performance. Meanwhile MAX-MIN is useful to maximize minimum performance, but fails to maximize overall performance in some situations. Future work will study the feasibility of the proposed solution through a detailed analysis of overheads and computations associated with the bandits’ operation.